# In celebration of Prof. Xun Yu Zhou’s 60th birthday

### v1

@@ 6 Existence of optimal control and convergence for general $\mu_{\rm ini}$}
$\leq$ $\int_{\mathbb{R}^{d}}|\log\rho^{\mu_{n}}(y)||f_{\bar{\mu}_{n}}(y)-f_{\bar{\mu}}(y)|dy+\int_{\mathbb{R}^{d}}|\log\rho^{\mu_{n}}(y)-\log\rho^{\mu}(y)|\bar{\mu}(dy)$
$\leq$ $\log\big(\frac{K}{\delta}\big)\|f_{\bar{\mu}_{n}}-f_{\bar{\mu}}\|_{\mathbb{L}^{1}}+\int_{\mathbb{R}^{d}}|\log\rho^{\mu_{n}}(y)-\log\rho^{\mu}(y)|\bar{\mu}(dy)\to 0,\quad\mbox{as $n\to\infty$.}$
Finally, since $G(\cdot)$ is continuous on $\mathscr{P}_{2}(\mathbb{R}^{d})$, by definition (6.9) we see that $F^{k}_{n}(\bar{\mu}_{n})\to F^{k}(\bar{\mu})$ whenever $\bar{\mu}_{n}\Rightarrow\bar{\mu}$. Thus by (6.12) we see that $\{F^{k}_{n}\}$ $\Gamma$-converges to $F^{k}$, as $n\to\infty$. This completes the proof. ∎
@@
(2) Note that the function $\phi$ in Assumption 3.2-(ii) satisfies $\phi(y)e^{\lambda|y|^{2}}\leq C$, for some $\lambda>0$ and that $\xi(y)\sim\lambda^{\prime}|y|^{2}$, as $|y|\to\infty$, Assumption 6.2-(ii) amounts to saying that $\phi$ and $\xi$ are compatible.
@@
We begin by denoting $\mu^{k}:=\mathbb{P}_{X_{T}^{\widehat{\alpha}^{k}}}$, $^{\mu_{\rm tar}}:=\mathbb{P}_{X_{T}^{\widehat{\alpha}}}$, and let $\rho^{\mu_{k}}=\Gamma_{1}(\mu_{k})$, $k\in\mathbb{N}$, $\rho^{\mu_{\rm tar}}=\Gamma_{1}({\mu_{\rm tar}})$, respectively, as we defined before. Next, applying (5.12), we have
$\begin{split}|\rho^{\mu_{k}}(y)-\rho^{\mu_{\rm tar}}(y)|&=|e^{\psi^{\mu_{k}}(y)}f_{\mu_{k}}(y)-e^{\psi^{{\mu_{\rm tar}}}(y)}f_{\mu_{\rm tar}}(y)|\\ &\leq e^{\psi^{{\mu_{\rm tar}}}(y)}|f_{\mu_{k}}(y)-f_{\mu_{\rm tar}}(y)|+f_{\mu_{k}}(y)|e^{\psi^{\mu_{k}}(y)}-e^{\psi^{{\mu_{\rm tar}}}(y)}|.\end{split}$ (6.15)
Let us now recall some facts from §4. First, note that the optimality of $\mu_{k}$ implies that $G(\mu_{k})\leq\frac{C}{k}$ (cf. (4.3)), for some generic constant $C>0$ independent of $k$, which we shall allow to vary from line to line below. Thus, by virtue of Assumption 3.2-(ii), we can write
$|f_{\mu_{k}}(y)-f_{\mu_{\rm tar}}(y)|\leq\frac{C}{k}\phi(y),\qquad y\in\mathbb{R}^{d},$
@@
To complete the proof, let us recall that optimal strategies are of the form $\widehat{\alpha}^{k}(t,x)=\nabla\log h^{k}(t,x)$, $k\in\mathbb{N}$, and $\widehat{\alpha}(t,x)=\nabla\log h(t,x)$, and $h^{k}(t,x)$ and $h(t,x)$ are the solutions to the respective PDEs:
$\begin{cases}\partial_{t}h^{k}(t,x)+\mathscr{L}_{t}h^{k}(t,x)=0;\\ h^{k}(T,x)=\rho^{\mu_{k}}(x).\end{cases}\quad\quad\begin{cases}\partial_{t}h(t,x)+\mathscr{L}_{t}h(t,x)=0;\\ h(T,x)=\rho^{\mu_{\rm tar}}(x),\end{cases}$ (6.17)
Furthermore, noting that $h^{k}(t,x)=\mathbb{E}[\rho^{\mu_{k}}(X^{t,x}_{T})]$, $h(t,x)=\mathbb{E}[\rho^{{\mu_{\rm tar}}}(X^{t,x}_{T})]$, and by the Bismut-Elworthy-Li formula we have
$\nabla h^{k}(t,x)=\mathbb{E}[\rho^{\mu_{k}}(X^{t,x}_{T})N^{t,x}_{T}];\quad\nabla h(t,x)=\mathbb{E}[\rho^{{\mu_{\rm tar}}}(X^{t,x}_{T})N^{t,x}_{T}].$
Thus, we have $|\nabla h(t,x)|\leq\mathbb{E}[\eta^{2}(X^{t,x}_{T})]^{\frac{1}{2}}\mathbb{E}[|N^{t,x}_{T}|]^{\frac{1}{2}}\leq\frac{C_{R}}{\sqrt{T-t}}$, whenever $(t,x)\in[0,T]\times B_{R}$, and a similar argument as in (4.9) and (4.11), together with the estimate (6.16), leads to that
$|\nabla h^{k}(t,x)-\nabla h(t,x)|\leq\frac{C_{R}}{k\sqrt{T-t}};\quad|h^{k}(t,x)-h(t,x)|\leq\frac{C_{R}}{k},\quad(t,x)\in[0,T]\times B_{R}.$ (6.18)
Finally, we note that by definition the function $h$ is positive everywhere (unless $\rho^{\mu_{\rm tar}}\equiv 0$), and being the classical solution to the PDE (6.17) it is continuous. Thus, given $R>0$, $h(t,x)\geq\delta_{R}>0$, for all $(t,x)\in[0,T]\times B_{R}$. Since (6.18) implies that $h^{k}$ converges to $h$ uniformly on compacts, thus it must hold that $h^{k}(t,x)\geq\delta_{R}/2$, for $(t,x)\in[0,T]\times B_{R}$, and $k$ large enough. We therefore conclude, similar to (4.1), that
$|\widehat{\alpha}^{k}(t,x)-\widehat{\alpha}(t,x)|$ $\leq\left|\frac{\nabla h^{k}(t,x)-\nabla h(t,x)}{h^{k}(t,x)}\right|+|\nabla h(t,x)|\left|\frac{h(t,x)-h^{k}(t,x)}{h^{k}(t,x)h(t,x)}\right|\leq\frac{C_{R}}{k\sqrt{T-t}},$

### v2

@@ 6 Existence of optimal control and convergence for general $\mu_{\rm ini}$}
$\leq$ $\int_{\mathbb{R}^{d}}|\log\rho^{\mu_{n}}(y)-\log\rho^{\mu}(y)|f_{\bar{\mu}_{n}}(y)(dy)+\int_{\mathbb{R}^{d}}|\log\rho^{\mu}(y)||f_{\bar{\mu}_{n}}(y)-f_{\bar{\mu}}(y)|dy=:I_{1}+I_{2},$
where $I_{1}$ and $I_{2}$ are the two integrals on the right side above. We now prove that both $I_{1}$ and $I_{2}$ converge to 0, as $n\to\infty$. To see this, first note that by (5.12) we have
$I_{1}$ $\leq$ $\int_{\mathbb{R}_{d}}|\psi^{\mu_{n}}(y)-\psi^{\mu}(y)|f_{\bar{\mu}_{n}(y)}dy+\int_{\mathbb{R}_{d}}|\log f_{\mu_{n}}(y)-\log f_{\mu}(y)|f_{\bar{\mu}_{n}(y)}dy$
$\leq$ $W_{2}(\mu_{n},\mu)\int_{\mathbb{R}_{d}}f_{\bar{\mu}_{n}(y)}dy+K\int_{\mathbb{R}_{d}}|\log f_{\mu_{n}}(y)-\log f_{\mu}(y)|g(y)dy=:I_{11}+I_{12},$
where the last inequality is due to (5.13) and Assumption 5.4. Since $\mu_{n}\Rightarrow\mu$, again by Lemma 5.6-(ii) and [65, Theorem 7.12], we have $W_{2}(\mu_{n},\mu)\to 0$. Hence $I_{11}\to 0$ as $n\to\infty$. Furthermore, by Assumption 5.4-(i) we see that, for $\mu\in\mathscr{E}$, one has $|\log f_{\mu}(y)|\leq C|y|^{2}$ for some constant $C>0$, that is, $|\log f_{\mu_{n}}(y)-\log f_{\mu}(y)|g(y)\leq 2C|y|^{2}g(y)$ for $y\in\mathbb{R}^{d}$. On the other hand, applying Lemma 5.6-(iii), we have $f_{{\mu}_{n}}\to f_{{\mu}}$ in $\mathbb{L}^{1}$. Thus, possibly along a subsequence (may assume itself, by virtue of Heine’s Lemma), it holds $\log f_{{\mu}_{n}}(y)\to\log f_{{\mu}}(y)$, a.e. $y\in\mathbb{R}^{d}$, thanks to the continuity of $\log$ function. Therefore, by Assumption 5.4 and Dominated Convergence Theorem, we see that $I_{12}\to 0$ as $n\to\infty$, whence $I_{1}\to 0$ as $n\to\infty$. Similarly, by (5.12), (5.15), and Assumption 5.4-(i), we have
$I_{2}$ $\leq$ $\int_{\mathbb{R}^{d}}|\psi^{\mu}(y)||f_{\bar{\mu}_{n}}(y)-f_{\bar{\mu}}(y)|dy+\int_{\mathbb{R}^{d}}|\log f_{\mu}(y)||f_{\bar{\mu}_{n}}(y)-f_{\bar{\mu}}(y)|dy$
$\leq$ $2C\int_{\mathbb{R}^{d}}|y|^{2}|f_{\bar{\mu}_{n}}(y)-f_{\bar{\mu}}(y)|dy\to 0,\quad\mbox{as $n\to\infty$},$
again thanks to Assumption 5.4-(ii) and the Dominated Convergence Theorem. Finally, since $G(\cdot)$ is continuous on $\mathscr{P}_{2}(\mathbb{R}^{d})$, by definition (6.6) we see that $F^{k}_{n}(\bar{\mu}_{n})\to F^{k}(\bar{\mu})$ whenever $\bar{\mu}_{n}\Rightarrow\bar{\mu}$. Thus, by (6.9), $\{F^{k}_{n}\}$ $\Gamma$-converges to $F^{k}$, as $n\to\infty$. The proof is now complete. ∎
@@
(2) Note that the function $\phi$ in Assumption 3.2-(ii) satisfies $\phi(y)e^{\lambda|y|^{2}}\leq C$, for some $c>0$ and $\xi(y)\sim c|y|^{2}$, as $|y|\to\infty$, Assumption 6.2-(ii) amounts to saying that $\phi$ and $\xi$ are compatible.
@@
We begin by denoting $\mu_{k}:=\mathbb{P}_{X_{T}^{\widehat{\alpha}^{k}}}$, $^{\mu_{\rm tar}}:=\mathbb{P}_{X_{T}^{\widehat{\alpha}}}$, and let $\rho^{\mu_{k}}=\Gamma_{1}(\mu_{k})$, $k\in\mathbb{N}$, $\rho^{\mu_{\rm tar}}=\Gamma_{1}({\mu_{\rm tar}})$, respectively, as we defined before. Next, applying (5.12), we have
$\begin{split}|\rho^{\mu_{k}}(y)-\rho^{\mu_{\rm tar}}(y)|&=|e^{\psi^{\mu_{k}}(y)}f_{\mu_{k}}(y)-e^{\psi^{{\mu_{\rm tar}}}(y)}f_{\rm tar}(y)|\\ &\leq e^{\psi^{{\mu_{\rm tar}}}(y)}|f_{\mu_{k}}(y)-f_{\rm tar}(y)|+f_{\mu_{k}}(y)|e^{\psi^{\mu_{k}}(y)}-e^{\psi^{{\mu_{\rm tar}}}(y)}|.\end{split}$ (6.12)
Let us now recall some facts from Section 4.1. First, note that the optimality of $\mu_{k}$ implies that $G(\mu_{k})\leq\frac{C}{k}$ (cf. (4.3)), for some generic constant $C>0$ independent of $k$, which we shall allow to vary from line to line below. Thus, by virtue of Assumption 3.2-(ii), we can write
$|f_{\mu_{k}}(y)-f_{\rm tar}(y)|\leq\frac{C}{k}\phi(y),\qquad y\in\mathbb{R}^{d},$
@@
To complete the proof, let us recall that optimal strategies are of the form $\widehat{\alpha}^{k}(t,x)=\nabla\log h^{\mu_{k}}(t,x)$, $k\in\mathbb{N}$, and $\widehat{\alpha}(t,x)=\nabla\log h(t,x)$, and $h^{\mu_{k}}(t,x)$ and $h(t,x)$ are the solutions to the respective PDEs:
$\begin{cases}\partial_{t}h^{\mu_{k}}(t,x)+\mathscr{L}_{t}h^{\mu_{k}}(t,x)=0;\\ h^{\mu_{k}}(T,x)=\rho^{\mu_{k}}(x).\end{cases}\quad\quad\begin{cases}\partial_{t}h(t,x)+\mathscr{L}_{t}h(t,x)=0;\\ h(T,x)=\rho^{\mu_{\rm tar}}(x),\end{cases}$ (6.14)
Furthermore, noting that $h^{\mu_{k}}(t,x)=\mathbb{E}[\rho^{\mu_{k}}(X^{t,x}_{T})]$, $h(t,x)=\mathbb{E}[\rho^{{\mu_{\rm tar}}}(X^{t,x}_{T})]$, and by the Bismut-Elworthy-Li formula we have
$\nabla h^{\mu_{k}}(t,x)=\mathbb{E}[\rho^{\mu_{k}}(X^{t,x}_{T})N^{t,x}_{T}];\quad\nabla h(t,x)=\mathbb{E}[\rho^{{\mu_{\rm tar}}}(X^{t,x}_{T})N^{t,x}_{T}].$
Thus, we have whenever $(t,x)\in[0,T]\times B_{R}$,
$|\nabla h(t,x)|\leq K\mathbb{E}[\eta^{2}(X^{t,x}_{T})]^{\frac{1}{2}}\mathbb{E}[|N^{t,x}_{T}|]^{\frac{1}{2}}\leq\frac{C_{R}}{\sqrt{T-t}},$
and a similar argument as in (4.8) and (4.10), together with the estimate (6.13), leads to that
$|\nabla h^{\mu_{k}}(t,x)-\nabla h(t,x)|\leq\frac{C_{R}}{k\sqrt{T-t}};\quad|h^{\mu_{k}}(t,x)-h(t,x)|\leq\frac{C_{R}}{k},\quad(t,x)\in[0,T]\times B_{R}.$ (6.15)
Finally, we note that by definition the function $h$ is positive everywhere, and being the classical solution to the PDE (6.14) it is continuous. Thus, given $R>0$, there exists $\delta_{R}$>0, such that $h(t,x)\geq\delta_{R}$, for all $(t,x)\in[0,T]\times B_{R}$. Since (6.15) implies that $h^{\mu_{k}}$ converges to $h$ uniformly on compacts, thus it must hold that $h^{\mu_{k}}(t,x)\geq\delta_{R}/2$, for $(t,x)\in[0,T]\times B_{R}$, and $k$ large enough. We therefore conclude, similar to (4.1), that
$|\widehat{\alpha}^{k}(t,x)-\widehat{\alpha}(t,x)|$ $\leq\left|\frac{\nabla h^{\mu_{k}}(t,x)-\nabla h(t,x)}{h^{\mu_{k}}(t,x)}\right|+|\nabla h(t,x)|\left|\frac{h(t,x)-h^{\mu_{k}}(t,x)}{h^{\mu_{k}}(t,x)h(t,x)}\right|\leq\frac{C_{R}}{k\sqrt{T-t}},$